4/x+3/5=10/x what is value of x
Answer by It is costly(175)

4/x+3/5=10/x
LCM=5x [4][5] +[3][x]=[10][5]
20+3x=50
20-20+3x=50-20
3x=30
3x/3=30/3
x=10